L^{-1} 1/s^{2}+4s+4 }


1
Expert's answer
2021-08-16T02:19:27-0400

"L^{-1}\left\{\frac{1}{s^2+4s+4}\right\}\\
\frac{1}{s^2+4s+4}=\frac{1}{\left(s+2\right)^2}\\
=L^{-1}\left\{\frac{1}{\left(s+2\right)^2}\right\}\\
\mathrm{Apply\:inverse\:transform\:rule:\quad if\:}L^{-1}\left\{F\left(s\right)\right\}=f\left(t\right)\mathrm{\:then}\:L^{-1}\left\{F\left(s-a\right)\right\}=e^{at}f\left(t\right)\\
\mathrm{For\:}\frac{1}{\left(s+2\right)^2}:\quad a=-2,\:\quad F\left(s\right)=\frac{1}{s^2}\\
=e^{-2t}L^{-1}\left\{\frac{1}{s^2}\right\}\\
=e^{-2t}t"